3. In a survey of a sample of 625 first-year college students, 371 of the students said that keeping up with political news is very important to them.

Our sample proportion would therefore be

A. 0.371. B. 0.460. C. 0.594. D. 0.625. E. 0.019
Answer #1

Solution

Given that,

x = 371

n=625

sampling proportion  =x / n = 371 / 625=0.594

answer = 0.594